Grade 11Trigonometry

If in a triangle ABC B = 60 degrees and sinA.sinC = ∂ . Find all possible values of ∂.

Kaustubh Nayyar
11 Years agoGrade 11

1 Answer

Ravi
ApprovedApproved Tutor Answer11 Years ago
Express C=120-A and solvesinA.sinC = ∂ to get the trigonometric solution of the eqn.

One such solution I got was a single trigonometric ratio , namely Cos(2A+60)=1/4-2∂

As -1<=Cos(2A+60)<=1

Using this inequality, you can get the values/ range of∂.